This is a four part question.
Part 1: Insert a number so the MEDIAN is 17:
15, 20, 18, _____
Part 2: Remove a number so that the mode is 45:
54, 63, 45, 76, 86, 45, 63, 81, 55 (Remove: _____)
Part 3: Remove a number so that the mean is 16.2:
9.9, 20, 6.5, 36, 12.1, 46.2, 12.7 (Remove: _____)
Part 4: What is the mean of all three answers?

Answer by stanbon(75887)

This is a four part question.
Part 1: Insert a number so the MEDIAN is 17:
15, 20, 18, _____
Sort the numbers: 15,18,20
Solve: (x+18)/2 = 17
x+18 = 34
x = 16
New sequence: 15,16,18,20
--------------------------------------------------------------
Part 2: Remove a number so that the mode is 45:
54, 63, 45, 76, 86, 45, 63, 81, 55 (Remove: _____)
Remove 63
-----------------------------------------------------------------
Part 3: Remove a number so that the mean is 16.2:
9.9, 20, 6.5, 36, 12.1, 46.2, 12.7 (Remove: _____)
Part 4: What is the mean of all three answers?
----
Current sum = 143.4
New sum after removal of one number = 6*16.2 = 97.2
--
Difference 143.4-97.2 = 46.2
Remove 46.2
-----------------------------------------
Cheers,
Stan H.
===============